# NoveltyBench: Evaluating Language Models for Humanlike Diversity

### v2

@@ Abstract
Language models have demonstrated remarkable capabilities on standard benchmarks, yet they struggle increasingly from mode collapse, the inability to generate diverse and novel outputs. Our work introduces NoveltyBench, a benchmark specifically designed to evaluate the ability of language models to produce multiple distinct and high-quality outputs. NoveltyBench utilizes prompts curated to elicit diverse answers and filtered real-world user queries. Evaluating 20 leading language models, we find that current state-of-the-art systems generate significantly less diversity than human writers. Notably, larger models within a family often exhibit less diversity than their smaller counterparts, challenging the notion that capability on standard benchmarks translates directly to generative utility. While prompting strategies like in-context regeneration can elicit diversity, our findings highlight a fundamental lack of distributional diversity in current models, reducing their utility for users seeking varied responses and suggesting the need for new training and evaluation paradigms that prioritize diversity alongside quality. [^1] Code, data, and more details are available at https://novelty-bench.github.io.
@@ 1 Introduction
Diversity of opinion, preference, and experience is a fundamental trait of being human. If you were to ask “Tell me a joke” or “What is the best book of all time?” to the next five people you talk to, with high likelihood you would receive five different answers. It is reasonable to expect language models to generate responses that have the same level of diversity as humans. Yet, when we ask models such as GPT-4 to recommend a movie or Claude 3 to suggest several vacation destinations, we often receive variations of the same few ideas—a phenomenon known as mode collapse.
@@ 3.1 Dataset Curation
We then had the authors of the paper write responses to each of the prompts, resulting in 8 human responses per prompt. Since this is a rather homogeneous set of human annotators (all work/study at the same institution), we expect this response set to represent an approximate lower bound on actual human diversity. However, as we will show in Section 4.3, even this group produced more diverse responses than most state-of-the-art language models.
@@ 3.2 Evaluating Novelty
Models with diversity alone would not be sufficient for users: each of the individual generations should also be of high quality, with the ultimate goal of maximizing the cumulative utility of a user who interacts with the model. Thus, our evaluation framework aims to assess the extent to which the set of responses is mutually diverse and simultaneously high-quality, and the our key metric directly models the cumulative utility for a user who observes a sequence of generations from a language model.
@@
How do we measure the cumulative utility of a sequence of generations $s_{1},s_{2},\ldots,s_{k}$? To achieve this, we first need a model of user behavior that describes how users interact with and consume language model generations. Inspired by the user model underlying the rank-biased precision metric in information retrieval (Moffat & Zobel, 2008), we assume that the user has a patience level $p\in[0,1]$: after observing each additional generation, they have a probability $p$ of requesting an additional generation from the language model and observing the next generation, and a probability $1-p$ of stopping interacting with the model. In other words, each additional generation should have a geometrically decaying effect on utility since it is more likely that the user simply would not have observed it.
@@ 4 Evaluation Results on NoveltyBench
Anthropic (Anthropic, 2024) Claude-3.5 Haiku 2.94 2.50
Claude-3.5 Sonnet 2.76 2.36
Claude-3 Opus 3.04 2.67
OpenAI (OpenAI, 2024) gpt-4o-mini 3.65 3.11
gpt-4o 3.88 3.27
Gemini (Google, 2024) gemini-1.5-pro 2.85 2.73
gemini-2.0-flash-lite 3.83 3.20
gemini-2.0-flash 3.81 3.17
gemini-2.0-pro 3.25 2.64
Cohere (Cohere, 2024) command-r7b 4.58 3.35
command-r 3.68 2.98
command-r-plus 3.79 3.08
Gemma 2 (Gemma Team et al., 2024) gemma-2-2b-it 6.66 4.63
gemma-2-9b-it 4.25 3.93
gemma-2-27b-it 4.03 3.77
Llama 3 (Llama Team et al., 2024) Llama-3.2-1B 7.74 2.81
Llama-3.2-3B 6.10 3.24
Llama-3.1-8B 6.24 3.76
Llama-3.3-70B 3.49 2.87
Llama-3.1-405B 4.20 3.39
@@ 4.2 State-of-the-art models struggle to generate diverse and useful outputs
Figure 2: Average number of unique generations out of a sample of 10 for all prompts in NoveltyBench.
Figure 3: Cumulative utility of generations of state-of-the-art models on NoveltyBench. A perfectly diverse and helpful model would have cumulative utility of 10.
Our evaluation reveals that frontier language models struggle to produce simultaneously diverse and useful responses. As shown in Figure 2, models like Claude 3 and GPT-4o produce on average fewer than 4 distinct responses in 10 queries. Notably, smaller models such as Gemma 2-2B and Llama 3.2-1B demonstrate the highest diversity on our benchmark. Figure 3 further illustrates this issue by measuring cumulative utility of generations. This metric captures both overall response diversity and individual generation quality. In our utility evaluation, closed-source models such as Claude 3, Gemini, and GPT-4o all scored below 4 out of a maximum utility score of 10, further indicating that even frontier models fall short of providing diverse and useful responses.
While larger models typically excel in standard benchmarks, they demonstrate worse diversity in their generations compared to smaller models. In Figure 4, we show how the cumulative utility of model generations changes as a function of the patience parameter in our evaluation setup. We focus exclusively on open-weight models (Gemma 2 and Llama 3), for which the number of parameters is publicly known. Again, at patience$=$0, we are essentially evaluating just a single generation regardless of the diversity of the output distribution. As patience increases, additional novel generations are given higher weights. Unsurprisingly, larger models perform better when only the best generation is considered. However, as users demand more diversity from the output distributions of these models, the larger models degrade in utility more quickly due to their lack of diversity compared to their smaller counterparts. Our results here suggest that progress on existing benchmarks that only consider a single output of the model can be misleading: larger models may appear superior when evaluated on single-generation tasks, but they can underperform smaller models when users require diverse, creative solutions across multiple interactions.
@@ Appendix A Additional Evaluation Details
A.1 Prompt for Selecting Diverse Prompts
@@
A.2 Example Prompts from NB-WildChat
@@
A.3 Evaluation of equivalence classification methods
To determine the equivalence of two language model generations, we used a variety of existing text similarity metrics such as BERTScore, BLEU, ROUGE, as well as fine-tuning OpenAI models (GPT-3.5-turbo and GPT-4o-mini) and a pre-trained DeBERTa model (microsoft/mdeberta-v3-large) on a manually labeled dataset of 1,000 pairs of LM-generated text.

### v4

@@ Abstract
Language models have demonstrated remarkable capabilities on standard benchmarks, yet they struggle increasingly from mode collapse, the inability to generate diverse and novel outputs. Our work introduces NoveltyBench, a benchmark specifically designed to evaluate the ability of language models to produce multiple distinct and high-quality outputs. NoveltyBench utilizes prompts curated to elicit diverse answers and filtered real-world user queries. Evaluating 20 leading language models, we find that current state-of-the-art systems generate significantly less diversity than human writers. Notably, larger models within a family often exhibit less diversity than their smaller counterparts, challenging the notion that capability on standard benchmarks translates directly to generative utility. While prompting strategies like in-context regeneration can elicit diversity, our findings highlight a fundamental lack of distributional diversity in current models, reducing their utility for users seeking varied responses and suggesting the need for new training and evaluation paradigms that prioritize diversity alongside quality. [^1] Code and data are available at https://novelty-bench.github.io.
@@ 1 Introduction
Diversity of opinion, preference, and experience is a fundamental trait of being human. If you were to ask “Tell me a joke” or “What is the best book of all time?” to the next five people you talk to, with high likelihood you would receive five different answers. It is reasonable to expect language models to generate responses that have the same level of diversity as humans. Yet, when we ask models such as GPT-4 to recommend a movie or Claude 3 to suggest several vacation destinations, we often receive variations of the same few ideas—a phenomenon known as mode collapse Hamilton (2024).
@@ 3.1 Dataset Curation
The authors of the paper then write responses to each of the prompts, resulting in 8 human responses per prompt. Since this is a rather homogeneous set of human annotators (all work/study at the same institution), we expect this response set to represent an approximate lower bound on actual human diversity. However, as we will show in Section 4.3, even this group produced more diverse responses than most state-of-the-art language models.
@@ 3.2 Evaluating Novelty
Models with diversity alone would not be sufficient for users: each of the individual generations should also be of high quality, with the ultimate goal of maximizing the cumulative utility of a user who interacts with the model. Thus, our evaluation framework aims to assess the extent to which the set of responses is mutually diverse and simultaneously high-quality, and our key metric directly models the cumulative utility for a user who observes a sequence of generations from a language model.
@@
How do we measure the cumulative utility of a sequence of generations $s_{1},s_{2},\ldots,s_{k}$? To achieve this, we first need a model of user behavior that describes how users interact with and consume language model generations. Inspired by the user model underlying the rank-biased precision metric in information retrieval (Moffat & Zobel, 2008), we assume that the user has a patience level $p\in[0,1]$: after observing each additional generation, they have a probability $p$ of requesting an additional generation from the language model and observing the next generation, and a probability 1p1-p of stopping interacting with the model. In other words, each additional generation should have a geometrically decaying effect on utility since it is more likely that the user simply would not have observed it.
@@ 4 Evaluation Results on NoveltyBench
Anthropic (Anthropic, 2024) Claude-3.5 Haiku 1.94 2.50
Claude-3.5 Sonnet 1.76 2.36
Claude-3 Opus 2.04 2.67
OpenAI (OpenAI, 2024) gpt-4o-mini 2.65 3.11
gpt-4o 2.88 3.27
Gemini (Google, 2024) gemini-1.5-pro 1.85 2.73
gemini-2.0-flash-lite 2.83 3.20
gemini-2.0-flash 2.81 3.17
gemini-2.0-pro 2.25 2.64
Cohere (Cohere, 2024) command-r7b 3.58 3.35
command-r 2.68 2.98
command-r-plus 2.79 3.08
Gemma 2 (Gemma Team et al., 2024) gemma-2-2b-it 5.66 4.63
gemma-2-9b-it 3.25 3.93
gemma-2-27b-it 3.03 3.77
Llama 3 (Llama Team et al., 2024) Llama-3.2-1B 6.74 2.81
Llama-3.2-3B 5.10 3.24
Llama-3.1-8B 5.24 3.76
Llama-3.3-70B 2.49 2.87
Llama-3.1-405B 3.20 3.39
@@ 4.2 State-of-the-art models struggle to generate diverse and useful outputs
Figure 2: Average number of unique generations out of a sample of 10 for all prompts in NoveltyBench. Error bars indicate 95% confidence intervals.
Figure 3: Cumulative utility of generations of state-of-the-art models on NoveltyBench. A perfectly diverse and helpful model would have cumulative utility of 10. Error bars indicate 95% confidence intervals.
Our evaluation reveals that frontier language models struggle to produce simultaneously diverse and useful responses. As shown in Figure 2, models like Claude 3 and GPT-4o produce on average fewer than 3 distinct responses in 10 queries. Notably, smaller models such as Gemma 2-2B and Llama 3.2-1B demonstrate the highest diversity on our benchmark. Figure 3 further illustrates this issue by measuring cumulative utility of generations. This metric captures both overall response diversity and individual generation quality. In our utility evaluation, closed-source models such as Claude 3, Gemini, and GPT-4o all scored below 4 out of a maximum utility score of 10, further indicating that even frontier models fall short of providing diverse and useful responses.
While larger models typically excel in standard benchmarks, they demonstrate worse diversity in their generations compared to smaller models. In Figure 4, we show how the cumulative utility of model generations changes as a function of the patience parameter in our evaluation setup. We focus exclusively on open-weight models (Gemma 2 and Llama 3), for which the number of parameters is publicly known. Again, at patience==0, we are essentially evaluating just a single generation regardless of the diversity of the output distribution. As patience increases, additional novel generations are given higher weights. Unsurprisingly, larger models perform better when only the best generation is considered. However, as users demand more diversity from the output distributions of these models, the larger models degrade in utility more quickly due to their lack of diversity compared to their smaller counterparts. Our results here suggest that progress on existing benchmarks that only consider a single output of the model can be misleading: larger models may appear superior when evaluated on single-generation tasks, but they can underperform smaller models when users require diverse, creative solutions across multiple interactions.
@@ 4.3 Alternate prompting methods for inducing diversity
4.4 How does alignment affect diversity?
The recent OLMo 2 series of models (OLMo et al., 2025), with weights released after each post-training stage, provides an opportunity to investigate the mechanistic effects of alignment on output diversity. We evaluate OLMo 2 models at all sizes (1B, 7B, 13B, 32B) after three stages of post-training: supervised fine-tuning (SFT), direct preference optimization (DPO), and finally reinforcement learning with verifiable rewards (RLVR) to analyze the effect of these three post-training stages on output diversity in Figure 6.
Figure 6: Diversity and utility analysis of OLMo 2 models across alignment stages.
Across model sizes, we observe that each alignment stage progressively reduces model diversity, with significant drops occurring during DPO. This finding suggests that current alignment procedures inadvertently suppress distributional diversity. Interestingly, the improvement in generation quality compensates for the loss in diversity. Specifically, we observe utility gains when transitioning from SFT to DPO, suggesting that while post-training algorithms may further reduce diversity, they can improve the overall usefulness of the remaining diverse outputs, highlighting a complex trade-off between alignment objectives and output diversity.
Our results suggest that alignment algorithms should be designed with output diversity in mind. Future work should investigate how to strike a better balance in alignment algorithms to achieve both safety and helpfulness objectives while maintaining the rich diversity that makes language models useful for creative tasks.
@@ Appendix A Additional Evaluation Details
A.1 Diversity of Reasoning Models
Figure 7: Diversity and utility performance of reasoning models on NoveltyBench. OpenAI’s models demonstrate a trend where increased reasoning effort leads to reduced diversity and utility.
Our core analysis focuses on the 20 most representative (non-reasoning) language models from major providers to establish a comprehensive baseline across model families and sizes, since reasoning capabilities are not yet widely available across model families at the time of writing. That said, as more frontier models carry out an explicit reasoning process before responding to user queries, we believe it would be interesting to see how such reasoning affects diversity.
To this end, we evaluate OpenAI and Google’s latest reasoning models (o3, o3-mini, and Gemini-2.5-flash) for which inference-time reasoning effort can be explicitly set, and explored how they impact diversity and utility on NoveltyBench. For the Gemini model, setting the reasoning budget had minimal effect on our evaluation. On the other hand, we see a trend with OpenAI’s models where additional reasoning seems to lead to less diverse and less useful outputs. We note that although reasoning models are usually used to solve problems with a single correct answer (e.g., math and coding), preserving diversity is still critical for the effectiveness of inference-time scaling strategies such as pass@k. Future work should further investigate how generation diversity affects creativity in problem solving, and its implications on inference-time scaling.
A.2 Prompt for Selecting Diverse Prompts
@@
A.3 Example Prompts from NB-WildChat
@@
A.4 Evaluation of equivalence classification methods
To determine the equivalence of two language model generations, we used a variety of existing text similarity metrics such as BERTScore, BLEU, ROUGE, as well as fine-tuning OpenAI models (GPT-3.5-turbo and GPT-4o-mini) and a pre-trained DeBERTa model (microsoft/deberta-v3-large) on a manually labeled dataset of 1,000 pairs of LM-generated text.